THEORETICAL RESULTS FOR (H_{2}O)_{n}H^{+} AND (H_{2}O)_{n}: QMC FOR H_{5}O_{2}^{+}, GRAPH THEORETICAL ENUMERATION OF DISTINCT HYDROGEN-BONDED ARRANGEMENTS, COMPACT TO EXTENDED PHASE TRANSITIONS IN (H_{2}O)_{n}H^{+}

\underline{SHERWIN J. SINGER}, LARS OJAMÄE and ISAIAH SHAVITT, Department of Chemistry, Ohio State University, Columbus, OH 43210.

Our recent progress in the study of protonated and neutral water clusters, (H_{2}O)_{n}H^{+} AND (H_{2}O)_{n} , is reviewed. Both variational and diffusion Monte Carlo calculations have been performed, leading to a prediction of the fundamental frequency for motion of the central proton in H_{5}O_{2}^{+}. Ab initio electronic structure calculations suggest this absorption feature will be intense. Extensive Monte Carlo simulations of larger hydrated proton clusters indicate there is a transition between compact and extended phases. Local minima of the potential energy surface for (H_{2}O)_{n}H^{+} AND (H_{2}O)_{n} are studied. The number of topologically distinct hydrogen-bonded arrangements can be enumerated using graph theory, leading to a prediction of the number of distinct local minima of the potential energy surface in compact geometries, including dodecahedral clathrate structures.